OpenStudy (anonymous):

surface area of small fig.=4ftsquared SA for big fig.= 'x' Volume of small fig.=8ft cubed Volume of big fig. =125ft cubed

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

If two 3D figures are similar, and x and y are the corresponding sides for the two figures, then Ratio of Sides: x/y Ratio of Areas: (x/y)^2 Ratio of Volumes: (x/y)^3

OpenStudy (anonymous):

but how do you set up the equation is it

OpenStudy (anonymous):

\[\frac{ 2 }{ x }=\frac{ 8 }{ 125 }\]

jimthompson5910 (jim_thompson5910):

I guess x and y was a bad choice, let me use a,b Ratio of Sides: a/b Ratio of Areas: (a/b)^2 Ratio of Volumes: (a/b)^3 ------------------------------------------------------- So if the surface areas are 4 and x, then we can say (a/b)^2 = 4/x which leads to a/b = sqrt(4/x) when you take the square root of both sides and if the volumes are 8 and 125, then (a/b)^3 = 8/125 ( sqrt(4/x) )^3 = 8/125 sqrt(4/x) = cube root of (8/125) sqrt(4/x) = 2/5 4/x = (2/5)^2 4/x = 4/25 x = ??

OpenStudy (anonymous):

ty ty <3

jimthompson5910 (jim_thompson5910):

np